OpenStudy (anonymous):

Solve this equation. 9^(5x)=27

OpenStudy (anonymous):

5x=3; x=3/5

OpenStudy (anonymous):

thats what i thought too but its incorrect

OpenStudy (anonymous):

Put in terms of 3: \[3^{2(5x)}=3^{3}\] Take log base 3 of both sides: \[\log_{3} (3^{10x})=\log_{3} 3^{3}\] \[10x=3\] \[x=3/10\]

OpenStudy (anonymous):

or just try writing both with base 3 and to hell with the log

OpenStudy (anonymous):

\[9^{5x}=3^{10x}\] \[27=3^3\] so \[3^{10x}=3^3\iff 10x=3\iff x=\frac{3}{10}\]